OpenStudy (anonymous):

Find exactly value sec(-11pi/12)

OpenStudy (anonymous):

sec(-11*180/12)=?

OpenStudy (anonymous):

\[\sec= (-\frac{ 11\pi }{ 12 })\]

OpenStudy (anonymous):

pi=180

OpenStudy (anonymous):

1/cos(-165)

OpenStudy (anonymous):

1st sec(-x) = secx

OpenStudy (anonymous):

sec(-11pi/12) = sec(11pie/12)